OpenStudy (anonymous):

The sum of first 16 terms of an A.P is 112 and sum of next 14 terms is 518. Find The A.P

OpenStudy (campbell_st):

ok... so the sum of an AP is \[S_{n} = \frac{n}{2}(2a + (n - 1)d)\] so you wull have 2 equations \[112 = \frac{16}{2}(2a + (16 -1) d)\] and \[518 = \frac{30}{2}(2a + (30 -1)d)\] when you simplify them you will have 2 equations in 2 unknowns... solve simultaneously.

OpenStudy (campbell_st):

can you clarify.... the sum of the next 14 terms is that 30 terms = 518 or terms 17 to 30... = 518